OpenStudy (anonymous):

.When p = 3, a = 15, and a varies directly with p, match each value of p with the corresponding value of a. TERM Definition p=5 A) a=20 p=6 B) a=25 p=7 C) a=30 p=4 D) a=35

jimthompson5910 (jim_thompson5910):

a = k*p 15 = k*3 15/3 = k 5 = k k = 5 ------------------------------------------------------- So the equation is a = 5p

OpenStudy (anonymous):

...

jimthompson5910 (jim_thompson5910):

So for instance, when p = 5... a = 5p a = 5*5 ... plug in p = 5 a = 25 which means p = 5 matches to a = 25

jimthompson5910 (jim_thompson5910):

you do the same for the rest

OpenStudy (anonymous):

i still dont get it

jimthompson5910 (jim_thompson5910):

when p = 6, what is 'a'?

jimthompson5910 (jim_thompson5910):

well plug it in a = 5p a = 5*6 a = ???

OpenStudy (anonymous):

i multiplied my answer with 5 :)

jimthompson5910 (jim_thompson5910):

to get ????

OpenStudy (anonymous):

what i did was 6*5 and got my answer

jimthompson5910 (jim_thompson5910):

and you got what

OpenStudy (anonymous):

30

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so p = 6 matches with a = 30

jimthompson5910 (jim_thompson5910):

and it sounds like you got the idea

OpenStudy (anonymous):

yes :)

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

i got to close this question to ask another....

jimthompson5910 (jim_thompson5910):

alright go ahead